# Imaging Shapes of Atomic Nuclei in High-Energy Nuclear Collisions

### v1

@@ Abstract
Values of observable in one collision system are heavily influenced by QGP properties during hydrodynamic evolution. These final state effects are largely mitigated by taking ratios between the two systems: $R_{\mathcal{O}}=\left\langle\mathcal{O}\right\rangle_{\mathrm{U+U}}/\left\langle\mathcal{O}\right\rangle_{\mathrm{Au+Au}}$ (Methods). Figure 3a–c show ratios for the three observables. Clearly, $R_{v_{2}^{2}}$ and $R_{(\delta p_{\mathrm{T}})^{2}}$ increase by up to 60% in central collisions, suggesting a large $\beta_{2\rm U}$, while $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$ decreases by up to threefold across centralities, indicating both a large $\beta_{2\rm U}$ and a small $\gamma_{\mathrm{U}}$. The ratios in the 0–5% most central range, having the greatest sensitivity to ${}^{238}$ shape, are visualized as hatch bands in Fig. 3d–f.
@@
Fig. 3: Constraining the shape of ${}^{238}$. Left column: ratios of $\left\langle v_{2}^{2}\right\rangle$ (a), $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ (b), and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ (c) between U+U and Au+Au collisions as a function of centrality. The data are compared to the IP-Glasma+MUSIC hydrodynamic model calculation assuming $\beta_{2\rm U}=0.28$ (red) and $\beta_{2\rm U}=0.25$ (blue), whose shaded bands denote the model uncertainties (Methods). Mid-column: ratio values in 0–5% most central collisions (hatch bands) for $\left\langle v_{2}^{2}\right\rangle$ (d) and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ (e), and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ (f) are compared to model calculations as a function of $\beta_{2\mathrm{U}}^{2}$ or $\beta_{2\mathrm{U}}^{3}$ for four $\gamma_{\mathrm{U}}$ values. The colored quadrilaterals delineates the allowed ranges of $\beta_{2\mathrm{U}}^{2}$ or $\beta_{2\mathrm{U}}^{3}$ from this data-model comparison. g shows the constrained ranges of ($\beta_{2\mathrm{U}}$,$\gamma_{\mathrm{U}}$) from three observables separately, and the confidence contours obtained by combining $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ (solid lines). The constraint from $R_{v_{2}^{2}}$ is viewed as a lower limit and hence is not used (see text).
@@
The data are compared with the state-of-the-art IP-Glasma+MUSIC hydrodynamic model Schenke et al. (2020a, b), which combines the fluctuating initial energy density distributions, relativistic viscous hydrodynamics, and hadronic transport. This model, successful in describing various flow observables at both RHIC and the LHC Schenke et al. (2020b), parameterizes nuclear shapes with a deformed Woods-Saxon profile,
@@
The deformation parameters for Au are fixed to an oblate shape with $\beta_{2\mathrm{Au}}=0.14$, while those for U are varied. The model also considers final state effects by adjusting QGP viscosities and initial condition uncertainties, including variations in nuclear radius $R_{0}$, skin $a$, $\beta_{2\mathrm{Au}}$, and higher-order shapes (Table 1 in Methods). These variations are included in the model uncertainties. The mass numbers of Au and U, differing by only 20%, result in almost complete cancellation of final state effects, leaving model uncertainties predominantly due to variations in initial conditions. In Fig. 3a-c, calculations for $\beta_{2\mathrm{U}}=0.28$ match central data for $R_{(\delta p_{\mathrm{T}})^{2}}$ and $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, but overestimates $R_{v_{2}^{2}}$ in 0–30% centrality. This overestimation could stem from the model’s limitations in describing $\varepsilon_{2}^{\mathrm{rp}}$-related $v_{2}$ components, which are more affected by structural parameters like nuclear radius and skin Jia et al. (2023). Thus, the $R_{v_{2}^{2}}$ comparison only sets a lower bound for $\beta_{2\mathrm{U}}$.
@@
In Fig. 3d-f, we contrast the 0–5% central data against predictions for varying $\beta_{2\mathrm{U}}$ and $\gamma_{\mathrm{U}}$. Calculated $R_{v_{2}^{2}}$ and $R_{(\delta p_{\mathrm{T}})^{2}}$ change linearly with $\beta_{2\mathrm{U}}^{2}$, while $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$ follows a $\beta_{2\mathrm{U}}^{3}\cos(3\gamma_{\mathrm{U}})$ trend, aligning remarkably with Eq. (2). The intersections between data and model delineate preferred $\beta_{2\mathrm{U}}$ ranges, yielding $\beta_{2\mathrm{U}}(R_{v_{2}^{2}})=0.240\pm 0.014$ and $\beta_{2\mathrm{U}}(R_{(\delta p_{\mathrm{T}})^{2}})=0.294\pm 0.021$. For $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, the favored $\beta_{2\mathrm{U}}$ range varies with $\gamma_{\mathrm{U}}$. These preferred ranges are summarized in Fig. 3g. Since $\beta_{2\mathrm{U}}(R_{v_{2}^{2}})$ value is viewed as a lower limit, a combined analysis of constraints from $R_{(\delta p_{\mathrm{T}})^{2}}$ and $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$ is performed, leading to $\beta_{2\mathrm{U}}=0.297\pm 0.013$ and $\gamma_{\mathrm{U}}=8.6^{\circ}\pm 4.8^{\circ}$ (mean and one standard deviation, see Methods).
@@
The extracted $\beta_{2\mathrm{U}}$ value is broadly in line with low-energy estimates Raman et al. (1987); Delaroche et al. (2010); Pritychenko et al. (2016), implying other sources of nucleon, quark and gluon correlations in ${}^{238}$ are less impactful compared to its large deformation, as supported by recent model studies Mäntysaari et al. (2023). Meanwhile, the nonzero $\gamma_{\mathrm{U}}$ value, highlighted by the confidence contours in Fig. 3g, suggests a small rigid triaxial component in ${}^{238}$ deformation, consistent with expectations from Ref. Otsuka et al. (2023). This marks the first indication of nuclear ground state triaxiality without involving transitions to excited states.
@@
Applications. We have demonstrated that flow-assisted imaging is a promising tool for exploring the structure of atomic nuclei in their ground state, complementary to the traditional approaches. The method’s strength lies in capturing a fast snapshot of nucleon spatial distribution, applicable to any collision species. This contrasts with nuclear spectroscopy, where complexity and interpretation vary with nucleus’s position on the nuclide chart. This approach is effective for discerning shape differences between species with similar mass numbers, ideally isobar pairs. Many applications are possible, with a few examples given here:
  • Odd-mass nuclides: For odd-mass nuclei, where either $N$ or $Z$ is odd, the nuclear shape is often presumed or determined via laser spectroscopy method Yang et al. (2023) to be similar to adjacent even-even nuclei. However, the spectroscopic data are more complex Bohr and Mottelson (1998) due to the coupling of the single unpaired nucleon with the nuclear core. By comparing the flow observables of odd-mass nuclei to selected even-even neighbors with known shapes, the high-energy approach avoids this complication,
@@
  • Dynamic deformations in “soft” nuclei: This method could distinguish between average deformation and transient shape fluctuations via measurements of multi-particle correlations Dimri et al. (2023); Acharya et al. (2021). This technique, sensitive only to the ground state, also sidesteps the complexities of disentangling shape variations during transitions to nuclear excited states.
  • $0\nu\beta\beta$ decay: The decay rate hinges on nuclear matrix elements (NME), significantly affected by the shapes of the initial and final species – a pair of isobars with the same mass number. Present NME uncertainties, partly stemming from inadequate knowledge of nuclear shapes, pose a major challenge in experimental design Engel and Menéndez (2017). This method, tailored for isobars, allows for the precise determination of shape differences between these species. This could reduce NME uncertainties, and hence aid in experiments searching for $0\nu\beta\beta$ decay and enhance our understanding of neutrino properties.
It would be remiss not to mention that this approach also holds promise in advancing the study of QGP, particularly its dynamics and transport properties, which have been limited by a poor understanding of QGP initial conditions Shuryak (2017); Busza et al. (2018) (Fig. 1e). By contrasting flow observables in similarly massed but structurally different species, our technique effectively eliminates final-state effects, thereby isolating initial condition variations seeded by shape differences. This can elucidate mechanisms of initial condition formation, consequently help to improve QGP transport property extraction through Bayesian inferences Bernhard et al. (2019); Everett et al. (2021); Nijs et al. (2021), and lead to breakthroughs in high-energy nuclear physics.
Collective flow assisted nuclear shape imaging is a discovery tool for exploring nuclear structure and high-energy nuclear collision physics. Future research could leverage colliders to conduct experiments with selected isobaric or isobar-like pairs. The combination of high- and low-energy techniques enables interdisciplinary opportunities in the study of atomic nuclei across energy scales.
@@ Methods
.1 Accessing information in the intrinsic frame
The nuclear shape in the intrinsic frame was considered not directly observable in low-energy experiments. In high-energy collisions, the collective flow phenomenon can be used to infer the shape and size of the nucleon distribution in the overlap region of the transverse plane, represented as $\rho\left(\mathbf{r}\right)$ with ${\bf r}=x+iy$, and hence connect directly to the shape characteristics of the two colliding nuclei in their intrinsic frame as discussed below.
The elliptic shape of the heavy-ion initial state is characterized by its amplitude $\varepsilon_{2}$ and direction $\Phi_{2}$, which are defined by nucleon positions as,
@@
It is important to note that when the coordinate system is rotated such that $x$ and $y$ coincide with the minor and major axes, the elliptic eccentricity coincides with the usual definition $\varepsilon_{2}=\frac{\left\langle y^{2}\right\rangle-\left\langle x^{2}\right\rangle}{\left\langle y^{2}\right\rangle+\left\langle x^{2}\right\rangle}$.
The parameter $\varepsilon_{2}$ plays a crucial role in driving the elliptic flow $v_{2}$ of the particle distribution: $v_{2}\propto\varepsilon_{2}$. In general, $n$-particle correlations quantify the $n^{\rm th}$-moments of elliptic flow and radial flow distributions, which are linearly related to the $n^{\rm th}$-moments of $\varepsilon_{2}$ and $d_{\perp}$.
To demonstrate this point, let’s consider collisions at zero impact parameter without loss of generality, where average elliptic geometry vanishes $\left\langle\mathcal{E}_{2}\right\rangle=0$. The second moment of eccentricity for many events becomes Giacalone et al. (2019); Giacalone (2023),
@@
where $\left\langle\rho(\mathbf{r})\right\rangle$ is the event-averaged profile, and
@@
is the usual two-body distribution. Similarly, the third central moments are related to the three-body distribution $\rho\left(\mathbf{r}_{1},\mathbf{r}_{2},\mathbf{r}_{3}\right)=\left\langle\delta\rho(\mathbf{r}_{1})\delta\rho(\mathbf{r}_{2})\delta\rho(\mathbf{r}_{3})\right\rangle$. For example,
$\left\langle\varepsilon_{2}^{2}\delta d_{\perp}/d_{\perp}\right\rangle$ $\approx-\frac{\int_{\mathbf{r}_{1},\mathbf{r}_{2},\mathbf{r}_{3}}\left(\mathbf{r}_{1}\right)^{2}\left(\mathbf{r}_{2}^{*}\right)^{2}|\mathbf{r}_{3}^{2}|\rho\left(\mathbf{r}_{1},\mathbf{r}_{2},\mathbf{r}_{3}\right)}{\left(\int_{\mathbf{r}}|\mathbf{r}|^{2}\left\langle\rho(\mathbf{r})\right\rangle\right)^{2}\int_{\mathbf{r}}|\mathbf{r}|^{2}\left\langle\rho(\mathbf{r})\right\rangle}\;,$ (6)
where we denote $\delta d_{\perp}/d_{\perp}\equiv(d_{\perp}-\left\langle d_{\perp}\right\rangle)/\left\langle d_{\perp}\right\rangle$, and the relation $\frac{\delta d_{\perp}}{d_{\perp}}\approx-\frac{\delta\left\langle|\mathbf{r}^{2}|\right\rangle}{\left\langle|\mathbf{r}^{2}|\right\rangle}=-\frac{\int_{\mathbf{r}}|\mathbf{r}^{2}|\delta\rho\left(\mathbf{r}\right)}{\int_{\mathbf{r}}|\mathbf{r}|^{2}\langle\rho(\mathbf{r})\rangle}$ is used.
@@
Here the scalar $\delta_{d}$ and vector $\mathcal{E}_{0}\equiv\varepsilon_{0}e^{2i\Phi_{2;0}}$ represent values for spherical nuclei. The values of scalar $p_{0}$ and vector ${\bm{p}}_{2}$ are directly connected to the $xy$-projected one-body distribution $\rho(\mathbf{r})$. Therefore, they depend on the orientation of the two nuclei. The fluctuations of $\delta_{d}$ ($\mathcal{E}_{0}$) are uncorrelated with $p_{0}$ (${\bm{p}}_{2}$). After averaging over collisions with different Euler angles and setting $\beta_{2p}=\beta_{2t}$ and $\gamma_{p}=\gamma_{t}$, we obtain:
@@
The event-averaged moments in Eq. (8) are rotationally invariant and capture the intrinsic many-body distributions of $\rho\left(\mathbf{r}\right)$. Note that the coefficients $a_{n}$ in Eq. (2) are strong functions of centrality that decrease towards central collisions, while coefficients $b_{n}$ change weakly with centrality. Therefore, the impact of deformation is always largest in the most central collisions. In general, it can be shown that the $n$-particle correlations reflect the rotational invariant ${n}^{\mathrm{th}}$ central moments of $\rho\left(\mathbf{r}\right)$, which in turn are connected to the ${n}^{\mathrm{th}}$ moments of the nuclear shape in the intrinsic frame.
.2 Previous experimental attempts on nuclear shapes at high energy
The idea that $v_{2}$ can be enhanced by $\beta_{2}$ was recognized early Li (2000); Shuryak (2000); Filip et al. (2009); Shou et al. (2015); Goldschmidt et al. (2015). Studies at RHIC Adamczyk et al. (2015a) and the LHC ALICE Collaboration (2018); Sirunyan et al. (2019); Aad et al. (2020) in ${}^{238}$+${}^{238}$ and ${}^{129}$e+${}^{129}$e collisions indicated $\beta_{2}$’s influence on $v_{2}$. Several theoretical investigations followed this up to assess the extent to which $\beta_{2}$ can be constrained by $v_{2}$ alone Giacalone et al. (2018); Giacalone (2020b); Giacalone et al. (2021); Ryssens et al. (2023). A challenge with $v_{2}$ is that its $a_{2}$ term in Eq. (2) is impacted by $\varepsilon_{2}^{\mathrm{rp}}$ and often exceeds the $b_{1}\beta_{2}^{2}$ even in central collisions. A recent measurement of $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ aimed to assess the triaxiality of ${}^{129}$e Aad et al. (2023), but the extraction of $\gamma_{\mathrm{Xe}}$ was hindered by needing prior knowledge of $\beta_{2\mathrm{Xe}}$ and potentially significant fluctuations in $\gamma_{\mathrm{Xe}}$ Robledo et al. (2008); Bally et al. (2022a, b); Dimri et al. (2023). The combination of multiple observables allows for more quantitative extraction of nuclear shape parameters, aided by Eq. (2).
.3 Event selection

### v2

@@ Abstract
Observables in one collision system are strongly influenced by QGP properties during hydrodynamic evolution. By taking ratios between the two systems, such final state effects are largely mitigated: $R_{\mathcal{O}}=\left\langle\mathcal{O}\right\rangle_{\mathrm{U+U}}/\left\langle\mathcal{O}\right\rangle_{\mathrm{Au+Au}}$ (Methods). Figure 3a–c show ratios for the three observables. $R_{v_{2}^{2}}$ and $R_{(\delta p_{\mathrm{T}})^{2}}$ increase by up to 60% in central collisions, requiring a large $\beta_{2\rm U}$, whereas $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$ decreases by up to threefold across centralities, demanding a large $\beta_{2\rm U}$ and a small $\gamma_{\mathrm{U}}$. The ratios in the 0–5% most central range, having the greatest sensitivity to 238U shape, are shown as hatch bands in Fig. 3d–f.
@@
Fig. 3: Constraining the shape of 238U. Left column: ratios of $\left\langle v_{2}^{2}\right\rangle$ (a), $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ (b), and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ (c) between U+U and Au+Au collisions as a function of centrality. The data are compared with the IP-Glasma+MUSIC hydrodynamic model calculation assuming $\beta_{2\rm U}=0.28$ (red) and $\beta_{2\rm U}=0.25$ (blue), shaded bands of which denote the model uncertainties (Methods). Mid-column: ratio values in 0–5% most central collisions (hatch bands) for $\left\langle v_{2}^{2}\right\rangle$ (d), $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ (e), and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ (f) are compared with model calculations as a function of $\beta_{2\mathrm{U}}^{2}$ or $\beta_{2\mathrm{U}}^{3}$ for four $\gamma_{\mathrm{U}}$ values. The colored quadrilaterals delineate the allowed ranges of $\beta_{2\mathrm{U}}^{2}$ or $\beta_{2\mathrm{U}}^{3}$ from this data-model comparison. g, the constrained ranges of ($\beta_{2\mathrm{U}}$,$\gamma_{\mathrm{U}}$) from three observables separately, and the confidence contours obtained by combining $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ (solid lines). The constraint from $R_{v_{2}^{2}}$ is viewed as a lower limit and hence is not used (see main text).
@@
The data are compared with the state-of-the-art IP-Glasma+MUSIC hydrodynamic model Schenke et al. (2020a, b), which combines the fluctuating initial energy density distributions, relativistic viscous hydrodynamics, and hadronic transport. This model, successful in describing flow observables at both RHIC and the LHC Schenke et al. (2020b), parameterizes nuclear shapes with a deformed Woods-Saxon profile,
@@
The parameters for Au are fixed to an oblate shape with $\beta_{2\mathrm{Au}}=0.14$ and $\gamma_{\mathrm{Au}}=45^{\circ}$ Bally et al. (2023), whereas those for U are varied. The model also considers final state effects by adjusting QGP viscosities and initial condition uncertainties, including variations in nuclear radius $R_{0}$, skin $a$, $\beta_{2\mathrm{Au}}$, $\gamma_{\mathrm{Au}}$, and higher-order shapes (Table 1 in Methods). These variations are included in the model uncertainties. The mass numbers of Au and U, differing by only 20%, result in almost completely canceling final state effects, leaving model uncertainties mainly from initial conditions. In Fig. 3a–c, calculations for $\beta_{2\mathrm{U}}=0.28$ match central data for $R_{(\delta p_{\mathrm{T}})^{2}}$ and $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, but overestimate $R_{v_{2}^{2}}$ in 0–30% centrality. This overestimation stems from the limitations of the model in describing $\varepsilon_{2}^{\mathrm{rp}}$-induced $v_{2}$ components, which are strongly affected by variations in the impact parameter and other structural parameters like nuclear radius and skin Jia et al. (2023), as well as possible longitudinal flow decorrelations Jia et al. (2024). Thus, the $R_{v_{2}^{2}}$ comparison is expected to set only a lower bound for $\beta_{2\mathrm{U}}$ in this model.
@@
In Fig. 3d-f, we contrast the 0–5% central data against predictions for varying $\beta_{2\mathrm{U}}$ and $\gamma_{\mathrm{U}}$. Calculated $R_{v_{2}^{2}}$ and $R_{(\delta p_{\mathrm{T}})^{2}}$ change linearly with $\beta_{2\mathrm{U}}^{2}$, whereas $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$ follows a $\beta_{2\mathrm{U}}^{3}\cos(3\gamma_{\mathrm{U}})$ trend, aligning remarkably with Eq. (2). The intersections between data and model delineate preferred $\beta_{2\mathrm{U}}$ ranges, yielding $\beta_{2\mathrm{U}}(R_{(\delta p_{\mathrm{T}})^{2}})=0.294\pm 0.021$ and a lower limit value $\beta_{2\mathrm{U}}(R_{v_{2}^{2}})=0.234\pm 0.014$. For $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, the favored $\beta_{2\mathrm{U}}$ range varies with $\gamma_{\mathrm{U}}$. These preferred ranges are shown in Fig. 3g. A combined analysis of constraints from $R_{(\delta p_{\mathrm{T}})^{2}}$ and $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$ is performed, yielding $\beta_{2\mathrm{U}}=0.297\pm 0.015$ and $\gamma_{\mathrm{U}}=8.5^{\circ}\pm 4.8^{\circ}$ (mean and one standard deviation, see Methods).
The data are also compared with Trajectum Giacalone et al. (2023), another hydrodynamic model with a different implementation of the initial condition and QGP evolution. Trajectum derives constraints on the initial and final state parameters of the QGP based on a Bayesian analysis of the LHC data that are then extrapolated to the RHIC energies. These constraints were not tuned to the RHIC data but are nevertheless useful in estimating the theoretical uncertainties. The relavent constraints are $\beta_{2\mathrm{U}}=0.275\pm 0.017$ and $\gamma_{\mathrm{U}}=15.5^{\circ}\pm 7.8^{\circ}$. A combination of constraints from the two models yields $\beta_{2\mathrm{U}}=0.286\pm 0.025$ and $\gamma_{\mathrm{U}}=8.7^{\circ}\pm 4.5^{\circ}$ (Methods).
@@
The extracted $\beta_{2\mathrm{U}}$ value is in line with low-energy estimates Pritychenko et al. (2016), implying other sources of nucleon, quark and gluon correlations in 238U are less impactful compared with its large deformation, as supported by recent model studies Mäntysaari et al. (2023). Meanwhile, the small $\gamma_{\mathrm{U}}$ value excludes a large triaxiality of Uranium and indicates an average value remarkably consistent with the low energy estimate based on a similar rigid-rotor assumption. This marks the first extraction of nuclear ground state triaxiality without involving transitions to excited states.
@@
Applications. The flow-assisted nuclear shape imaging is a promising tool for exploring the structure of atomic nuclei in their ground state. The strength of this method lies in capturing a fast snapshot of nucleon spatial distribution, applicable to any collision species. This contrasts with nuclear spectroscopy, in which complexity varies with the position of nucleus on the Segrè chart. This approach is effective for discerning shape differences between species with similar mass numbers, ideally isobar pairs. Many applications are possible, with a few examples given here:
  • Odd-mass nuclides: for odd-mass nuclei, where either $N$ or $Z$ is odd, the nuclear shapes should be similar to adjacent even-even nuclei. As the transition data are more complex Bohr and Mottelson (1998), the ground-state shapes are usually inferred from the data measured via laser spectroscopy method Yang et al. (2023). The high-energy approach is suitable for both odd-mass nuclei and even-even nuclei, hence eliminating a possible source of bias in low-energy experiments.
@@
  • Dynamic deformations in “soft” nuclei: this method could distinguish between average deformation and transient shape fluctuations via measurements of multi-particle correlations Dimri et al. (2023). For example, the sixth-particle correlator $\left\langle v_{2}^{6}\right\rangle$ has direct sensitivity to the fluctuations of $\gamma$. This information was obtained in rare cases at low energy Wu et al. (1996); Ayangeakaa et al. (2019). Our technique, sensitive only to the ground state, also sidesteps the complexities of disentangling shape variations during transitions to excited states.
  • $0\nu\beta\beta$ decay: the decay rate hinges on nuclear matrix elements (NME), significantly affected by the shapes of the initial and final species – a pair of isobars with the same mass number. Present NME uncertainties, partly stemming from inadequate knowledge of nuclear shapes, pose a main challenge in experimental design Engel and Menéndez (2017). This method, tailored for isobars, allows for precisely determining shape differences between these species. This could reduce NME uncertainties, and hence aid in experiments searching for $0\nu\beta\beta$ decay and enhance our understanding of neutrino properties.
It would be remiss not to mention that our approach also holds promise in advancing the study of QGP, particularly its dynamics and transport properties, which have been limited by a poor understanding of QGP initial conditions Shuryak (2017); Busza et al. (2018) (Fig. 1e). By contrasting flow observables in similarly massed but structurally different species, our technique effectively eliminates final-state effects, thereby isolating initial condition variations seeded by shape differences. This can explain the mechanisms of initial condition formation and consequently help to improve QGP transport property extraction through Bayesian inferences Bernhard et al. (2019); Everett et al. (2021); Nijs et al. (2021) and lead to breakthroughs in high-energy nuclear physics.
Collective flow assisted nuclear shape imaging is a discovery tool for exploring nuclear structure and high-energy nuclear collision physics. Future research could leverage colliders to conduct experiments with selected isobaric or isobar-like pairs. The combination of high- and low-energy techniques enables interdisciplinary research in the study of atomic nuclei across energy scales.
.1 Online content
Any methods, additional references, Nature Portfolio reporting summaries, source data, extended data, supplementary information, acknowledgements, peer review information; details of author contributions and competing interests; and statements of data and code availability are available at https://doi.org/10.1038/s41586-024-08097-2 .
@@ Methods
.2 Accessing information in the intrinsic frame
The nuclear shape in the intrinsic frame is not directly observable in low-energy experiments. However, in high-energy collisions, the collective flow phenomenon is sensitive to the shape and size of the nucleon distribution in the overlap region of the transverse plane. This distribution, denoted as $\rho\left(\mathbf{r}\right)$ with ${\bf r}=x+iy$, provides a direct link to the shape characteristics of the two colliding nuclei in their intrinsic frames, as discussed below.
The elliptic shape of the heavy-ion initial state is characterized by its amplitude $\varepsilon_{2}$ and direction $\Phi_{2}$, defined by nucleon positions as,
@@
When the coordinate system is rotated such that $x$ and $y$ coincide with the minor and major axes, the elliptic eccentricity coincides with the usual definition $\varepsilon_{2}=\frac{\left\langle y^{2}\right\rangle-\left\langle x^{2}\right\rangle}{\left\langle y^{2}\right\rangle+\left\langle x^{2}\right\rangle}$. The parameter $\varepsilon_{2}$ drives the elliptic flow $v_{2}$: $v_{2}\propto\varepsilon_{2}$.
Let’s now consider collisions at zero impact parameter, where, without loss of generality, the average elliptic geometry vanishes, i.e., $\left\langle\mathcal{E}_{2}\right\rangle=0$. The second moment of eccentricity over many events is given by Giacalone et al. (2019); Giacalone (2023),
@@
where $\left\langle\rho(\mathbf{r})\right\rangle$ represents the event-averaged profile, and
@@
is the usual two-body distribution. Similarly, the third central moments are related to the three-body distribution, $\rho\left(\mathbf{r}_{1},\mathbf{r}_{2},\mathbf{r}_{3}\right)=\left\langle\delta\rho(\mathbf{r}_{1})\delta\rho(\mathbf{r}_{2})\delta\rho(\mathbf{r}_{3})\right\rangle$. For example,
$\left\langle\varepsilon_{2}^{2}\delta d_{\perp}/d_{\perp}\right\rangle$ $\approx-\frac{\int_{\mathbf{r}_{1},\mathbf{r}_{2},\mathbf{r}_{3}}\left(\mathbf{r}_{1}\right)^{2}\left(\mathbf{r}_{2}^{*}\right)^{2}|\mathbf{r}_{3}^{2}|\rho\left(\mathbf{r}_{1},\mathbf{r}_{2},\mathbf{r}_{3}\right)}{\left(\int_{\mathbf{r}}|\mathbf{r}|^{2}\left\langle\rho(\mathbf{r})\right\rangle\right)^{3}}\;,$ (6)
where we define $\delta d_{\perp}/d_{\perp}\equiv(d_{\perp}-\left\langle d_{\perp}\right\rangle)/\left\langle d_{\perp}\right\rangle$, and the relation $\frac{\delta d_{\perp}}{d_{\perp}}\approx-\frac{\delta\left\langle|\mathbf{r}^{2}|\right\rangle}{\left\langle|\mathbf{r}^{2}|\right\rangle}=-\frac{\int_{\mathbf{r}}|\mathbf{r}^{2}|\delta\rho\left(\mathbf{r}\right)}{\int_{\mathbf{r}}|\mathbf{r}|^{2}\langle\rho(\mathbf{r})\rangle}$ is used.
@@
Here, the scalar $\delta_{d}$ and vector $\mathcal{E}_{0}$ represent values for spherical nuclei. The values of scalar $p_{0}$ and vector ${\bm{p}}_{2}$ are directly connected to the $xy$-projected one-body distribution $\rho(\mathbf{r})$. Therefore, they depend on the orientation of the two nuclei. The fluctuations of $\delta_{d}$ ($\mathcal{E}_{0}$) are uncorrelated with $p_{0}$ (${\bm{p}}_{2}$). After averaging over collisions with different Euler angles and setting $\beta_{2p}=\beta_{2t}$ and $\gamma_{p}=\gamma_{t}$, we obtain:
@@
The event-averaged moments in Eq. (8) are rotationally invariant and capture the intrinsic many-body distributions of $\rho\left(\mathbf{r}\right)$. Note that the coefficients $a_{n}$ in Eq. (2) are strong functions of centrality that decrease towards central collisions, while coefficients $b_{n}$ vary weakly with centrality. Therefore, the impact of deformation is always largest in the most central collisions. In general, it can be shown that the $n$-particle correlations reflect the rotational invariant ${n}^{\mathrm{th}}$ central moments of $\rho\left(\mathbf{r}\right)$, which in turn are connected to the ${n}^{\mathrm{th}}$ moments of the nuclear shape in the intrinsic frame.
.3 Previous experimental attempts on nuclear shapes at high energy
The idea that $v_{2}$ can be enhanced by $\beta_{2}$ was recognized early Li (2000); Shuryak (2000); Filip et al. (2009); Shou et al. (2015); Goldschmidt et al. (2015). Studies at RHIC Ada (2015) and the LHC ALICE Collaboration (2018); CMS Collaboration (2019); ATLAS Collaboration (2020) in 238U+238U and 129Xe+129Xe collisions indicated $\beta_{2}$’s influence on $v_{2}$. Several later theoretical investigations assessed the extent to which $\beta_{2}$ can be constrained by $v_{2}$ alone Giacalone et al. (2018); Giacalone (2020b); Giacalone et al. (2021); Ryssens et al. (2023). A challenge with $v_{2}$ is that its $a_{1}$ term in Eq. (2) is impacted by $\varepsilon_{2}^{\mathrm{rp}}$, which often exceeds the $b_{1}\beta_{2}^{2}$ term even in central collisions. A recent measurement of $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ aimed to assess the triaxiality of 129Xe Aad et al. (2023), but the extraction of $\gamma_{\mathrm{Xe}}$ was hindered by needing prior knowledge of $\beta_{2\mathrm{Xe}}$ and potentially substantial fluctuations in $\gamma_{\mathrm{Xe}}$ Robledo et al. (2008); Bally et al. (2022a, b); Dimri et al. (2023); Fortier et al. (2023). The combination of several observables in this article allows for more quantitative extraction of nuclear shape parameters.
.4 Event selection